OpenStudy (anonymous):

Solve the system of equations by the addition method. If the system does not have a single ordered pair as a solution, state whether the system is inconsistent or dependent. -4x - 3y = 6 -20x - 15y = 30 A. No solution, inconsistent B. (-24, -18) C. (-18, -24) D. Infinitely many solutions, dependent

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (texaschic101):

The addition method is also called the elimination method. Lets begin... -4x - 3y = 6 -->(-5) -20x - 15y = 30 ---------------- 20x + 15y = -30 (result of multiplying by -5) -20x - 15y = 30 --------------- 0 = 0 we do not need to go any further.....when you get an answer at the end that is equal, like 2 - 2 or 13 = 13, all that means is that the equations have infinite solutions. Now if it would have come out like 2 = 3 or 5 = 5, then there would have been no solutions or infinite solutions. Any questions ?

OpenStudy (texaschic101):

oops...that is supposed to be 2 = 2

OpenStudy (texaschic101):

oops...another typo....2 = 3 or 5 = -5, then there would have been no solutions or 0 solutions. sorry my brain is just waking up....lol

OpenStudy (anonymous):

mine keeps going back to sleep in the middle of class